# Logical Distillation of Graph Neural Networks

### v2

@@ 1 Introduction
We test IDTs on multiple synthetic and real-world datasets, performing distillation on two prominent GNN architectures, Graph Isomorphism Networks (GIN) (Xu et al., 2019) and Graph Convolution Networks (GCN) (Kipf and Welling, 2017). Our algorithm consistently distills IDTs that are succinct and have comparable predictive performance to the underlying GNN. Furthermore, when the ground truth is a $\mathrm{C}^{2}$ formula, the distilled IDT exhibits better generalization, outperforming the GNN on the test data. Qualitatively, we find that our method can provide new insights. For instance, on the AIDS dataset (Riesen and Bunke, 2008), IDTs infer a very simple high-performing rule that achieves over 99% classification accuracy. This rule classifies graphs based on their number of nodes being smaller or larger than 12. To the best of our knowledge, none of the existing GNNs or explanation methods have been able to infer this rule.
In the next section we discuss the relevant related work. In Section 3 we discuss the necessary background on graphs, logic and GNNs. We present IDTs in Section 4 and show how IDTs can be learned from GNNs in Section 5. In Section 6, we introduce an extension of $\mathrm{C}^{2}$, which allows us to learn more expressive IDTs. Finally, we empirically evaluate IDTs in Section 7. We analyze some of the obtained logical explanations in Section 7.4. We summarize our work and discuss future research directions in Section 8.
@@ 2 Related Work
Explanation methods that distill surrogate models from GNNs come closest to our approach. Azzolin et al. (2023) first derive instance-level local subgraphs as explanations and then cluster them to extrapolate a model-level Boolean formula using the subgraphs as concepts. Yuan et al. (2020) base their approach on input-optimization, i.e. globally reducing graphs to a number of instances for which the explanation is then given. Their approach requires prior domain knowledge. Most recently, Müller et al. (2024) first compute (almost) categorical layer wise node representations using GNN layers with Gumbel-Softmax update functions. Subsequently, they replace the neural networks by decision trees trained on the categorical node representations. Their approach results in an interpretable message passing scheme based on intermediate categorical node states, but requires to train a specific GNN architecture. All three approaches use graphs, sub-graphs or their combinations as the explanation model. This restricts these methods, as many simple and important constraints, e.g. a graph has more than 12 nodes, can not easily be expressed in terms of subgraphs.
@@ 3.3 Decision Trees
where $\text{path}(v)$ is the path from the root to $v$ and for each leaf set $M$ we define
@@ 4 Iterated Decision Trees
$\chi_{{\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}M_{2}}}\Longleftrightarrow$$\>\neg(AU_{1}>0)\vee((AU_{1}>0)\wedge(AU_{1}>1))$
$\chi_{{\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}M_{3}}}\Longleftrightarrow$$\>\neg(AU_{1}>0)\wedge\neg(AU_{1}>1)$
@@
$\chi_{{\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}M_{2}}}\Longleftrightarrow$$\>\neg(AU_{1}=1)$
$\chi_{{\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}M_{3}}}\Longleftrightarrow$ $\>(AU_{1}=1)$
@@
Consider an iterated decision tree consisting of two layers, the first layer as in Example 5. Now we can add a second layer that checks if there is more than one neighbor for which $\chi_{{\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}M_{2}}}=\chi_{{\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}M_{2}^{0}}}$ is true.
@@ 5 Learning Iterated Decision Trees
Recall the graph from Example 2. Assume that $\mathcal{G}$ consists of only this graph and graph label. Suppose that $\mathsf{LearnIDTLayer}(\mathcal{G},\mathcal{U},\mathcal{X}_{0})$ yields the IDT layer shown in Example 5. Then we extend the node attributes with binary vectors representing $\chi_{M}$ for each leaf set $M\in\{{\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}M_{0}},{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}M_{1}},{\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}M_{2}},{\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}M_{3}}\}$. For ${\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}M_{0}}$ we add
@@
For ${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}M_{1}},{\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}M_{2}},{\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}M_{3}}$ we add
@@
$\mathcal{U}=\begin{pmatrix}0&1&\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}1&\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}0&\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}1&\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}0\\ 1&0&\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}0&\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}1&\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}1&\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}0\\ 0&0&\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}0&\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}0&\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}0&\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}1\\ 1&1&\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}1&\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}0&\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}1&\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}0\end{pmatrix}$
@@ 5.2 Practical Considerations
    For the final iterated decision tree layer, we only consider the modal parameter $1$, corresponding to sum pooling.
@@
    We limit the depth of all but the final iterated decision tree layers to two to enhance interpretability. However, for the final IDT layer, we don’t limit the depth but instead perform minimal cost complexity pruning (Breiman et al., 1984).
@@ 8 Conclusion
We have presented Iterated Decision Trees (IDTs), a novel model tailored towards distilling interpretable logical formulas from Graph Neural Networks (GNNs). IDTs can express any logical formula expressible in first order logic with two variables and counting quantifiers ($\mathrm{C}^{2}$) — a fragment of first-order logic that is closely connected to the logical expressivity of GNNs. We have also introduced an extension of $\mathrm{C}^{2}$ that captures operations commonly used in GNNs. This extension was easily incorporated into IDTs without any significant computational overhead. The distilled IDTs often surpass the accuracy of the underlying GNN while providing insight into the decision process. They also outperform the considered GNNs when the ground truth is itself a logical formula. The classification decisions of the IDT are interpretable and enable us to extract insights on multiple datasets.

### v3

@@ 1 Introduction
We test IDTs on multiple synthetic and real-world datasets, performing distillation on two prominent GNN architectures: Graph Isomorphism Networks (GIN) (Xu et al., 2019) and Graph Convolution Networks (GCN) (Kipf and Welling, 2017). Our algorithm consistently distills IDTs that are succinct and have comparable predictive performance to the underlying GNN. Furthermore, when the ground truth is a $\mathrm{C}^{2}$ formula, the distilled IDT exhibits better generalization, outperforming the GNN on the test data. Qualitatively, we find that our method can provide new insights. For instance, on the AIDS dataset (Riesen and Bunke, 2008), IDTs infer a very simple high-performing rule that achieves over 99% classification accuracy. This rule classifies graphs based on their number of nodes being smaller or larger than 12. To the best of our knowledge, none of the existing GNNs or explanation methods have been able to infer this rule.
In the next section we discuss the relevant related work. In Section 3 we discuss the necessary background on graphs, logic and GNNs. We present IDTs in Section 4 and show how IDTs can be distilled from GNNs in Section 5. In Section 6, we introduce an extension of $\mathrm{C}^{2}$, which allows us to learn more expressive IDTs. We empirically evaluate IDTs in Section 7 and analyze some of the obtained logical explanations in Section 7.4. Finally, we summarize our work and discuss future research directions in Section 8.
@@ 2 Related Work
Explanation methods that distill surrogate models from GNNs come closest to our approach. Azzolin et al. (2023) first derive instance-level local subgraphs as explanations and then cluster them to extrapolate a model-level Boolean formula using the subgraphs as concepts. Yuan et al. (2020) base their approach on input-optimization, i.e. globally reducing graphs to a number of instances for which the explanation is then given. Their approach requires prior domain knowledge. Müller et al. (2024) first compute (almost) categorical layer wise node representations using GNN layers with Gumbel-Softmax update functions. Subsequently, they replace the neural networks by decision trees trained on the categorical node representations. Their approach results in an interpretable message passing scheme based on intermediate categorical node states but requires to train a specific GNN architecture. All three approaches use graphs, sub-graphs, or their combinations as the explanation model. This restricts these methods, as many simple and important constraints, e.g. a graph has more than 12 nodes, can not easily be expressed in terms of subgraphs. Similarly to us, Köhler and Heindorf (2024) use a description logic based method for explaining GNNs. Their method uses GNNs as a black box, whereas IDTs are distilled layer-by-layer, exploiting GNNs representations at each iteration. Furthermore, their approach currently does not support counting quantifiers.
@@ 3.3 Decision Trees
where $\text{path}(v)$ is the path from the root to $v$. For each set of leaves $M$ we define
@@ 4 Iterated Decision Trees
$\chi_{{\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}M_{2}}}\Longleftrightarrow$$\>\neg(AU_{1}>0)\vee((AU_{1}>0)\wedge(AU_{1}>1))$
$\chi_{{\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}M_{3}}}\Longleftrightarrow$$\>\neg(AU_{1}>0)\wedge\neg(AU_{1}>1)$
@@
$\chi_{{\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}M_{2}}}\Longleftrightarrow$$\>\neg(AU_{1}=1)$
$\chi_{{\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}M_{3}}}\Longleftrightarrow$ $\>(AU_{1}=1)$
@@
Consider an iterated decision tree consisting of two layers, the first layer as in Example 5. Now we can add a second layer that checks if there is more than one neighbor for which $\chi_{{\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}M_{2}}}=\chi_{{\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}M_{2}^{0}}}$ is true.
@@ 5 Learning Iterated Decision Trees
Recall the graph from Example 2. Assume that $\mathcal{G}$ consists of only this graph and graph label. Suppose that $\mathsf{LearnIDTLayer}(\mathcal{G},\mathcal{U},\mathcal{X}_{0})$ yields the IDT layer shown in Example 5. Then we extend the node attributes with binary vectors representing $\chi_{M}$ for each leaf set $M\in\{{\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}M_{0}},{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}M_{1}},{\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}M_{2}},{\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}M_{3}}\}$. For ${\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}M_{0}}$ we add
@@
For ${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}M_{1}},{\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}M_{2}},{\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}M_{3}}$ we add
@@
$\mathcal{U}=\begin{pmatrix}0&1&\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}1&\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}0&\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}1&\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}0\\ 1&0&\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}0&\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}1&\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}1&\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}0\\ 0&0&\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}0&\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}0&\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}0&\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}1\\ 1&1&\color[rgb]{0.84765625,0.10546875,0.375}\definecolor[named]{pgfstrokecolor}{rgb}{0.84765625,0.10546875,0.375}1&\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}0&\color[rgb]{0,0.30078125,0.25}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.30078125,0.25}1&\color[rgb]{1,0.7578125,0.02734375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7578125,0.02734375}0\end{pmatrix}$
@@ 5.2 Practical Considerations
    As our experiments consist of graph classification tasks, we only consider the modal parameter $1$ for the final iterated decision tree layer. A formula with an outer-most modal parameter $1$ provides identical labels for all nodes. Such a formula can thus be viewed as a graph classifier.
@@
    We limit the depth of all but the final iterated decision tree layers to two to enhance interpretability. For the final IDT layer, we don’t limit the depth but instead perform minimal cost complexity pruning (Breiman et al., 1984).
@@ 8 Conclusion
We have presented Iterated Decision Trees (IDTs), a novel model tailored towards distilling interpretable logical formulas from Graph Neural Networks (GNNs). IDTs can express any logical formula expressible in first order logic with two variables and counting quantifiers ($\mathrm{C}^{2}$) — a fragment of first-order logic that is closely connected to the logical expressivity of GNNs. We have also introduced an extension of $\mathrm{C}^{2}$ that captures operations commonly used in GNNs, such as mean aggregation. This extension was easily incorporated into IDTs without any significant computational overhead. The distilled IDTs often surpass the accuracy of the underlying GNN while providing insight into the decision process. They also outperform the considered GNNs when the ground truth is itself a logical formula. The classification decisions of the IDT are interpretable and enable us to extract insights on multiple datasets.
9 Future Work and Limitations
@@
Acknowledgments
This work was funded in part by the Vienna Science and Technology Fund (WWTF), project StruDL (ICT22-059); by the Austrian Science Fund (FWF), project NanOX-ML (6728); and by the European Unions Horizon Europe Doctoral Network programme under the Marie-Skłodowska-Curie grant, project Training Alliance for Computational systems chemistry (101072930). SM would like to thank Steve Azzolin for the fruitful discussions.